A 20 kg Emperor penguin consumes 4 kg of cold fish (0°C) on a foraging trip. The penguin's normal metabolic rate is 40 Watts. Assume the normal body temperature of a penguin is 40°C.

What proportion (expressed as a percentage value) of its energy budget over a 24 hour period (i.e. total Joules used in 24 hours to serve resting metabolism) would be required to warm the meal it just consumed up to penguin body temperature?
Answer #1

Specific heat capacity of the fish is about 3.6 kJ/kg K

Hence,

amount of heat required to warm up the fish

= m*c*(T2-T1)

= 4*3.6*(40-0)

= 576 kJ.

Normal metabolic rate is 40 W.

Hence, amount of energy generated in 24 hours

= 40*(24*60*60)

= 3456,000 J or 3456 kJ

Hence, proportion of energy used

= 576/3456

= 0.167 or 16.7%

Please leave a thumbs up if you like the answer!